INAF queue observing, 2014 Feb 04-05

Observer: E. Sani, F. Cusano, S. Bisogni, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn on call
Instrument: MODS

Summary

At twilight we thin cirrus pasassing from SE to NE, 2m/s blowing wind, no gusts and 20% humidity.
We have clouds passing, these become thick at ~8.30UT avoiding guiding and collimation.
The TO closes the dome at 9.30 due to clouds and increasing humidity.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard G191b2b dual grating - airmass 1.1
Vanzella_B macsj0647 dual grating 3h some clouds in the last 0.5h
Gilli_Q1 J1048 red grating 1h some clouds in the last 15min. MASK EXECTUTED

Details

00:45 The TO opens the dome

01:55 Slew to the standard G191b2b
  • acqMODS g191b2b__recovery.acq mods1r.201400205.0002-0003
  • since the acquisition for this Std was already done during 20140130UT, we run the recovery acquisition using the mods_lastOffset script (the procedure is described in the SciOpsLBTO webpages --> MODS --> Observing --> Recovery of Observations)

02:04 execMods g191b2b_ls.obs mods1b 0002 mods1r 0004
  • the average seeing on the GCS is 1.3

02:12 We slew to the scientific target Vanzella_20B MACS J0647
  • acqMODS mods.macsj0647.acq mods1r 0005-0009
  • modsAlign mods.macsj0647.mms mods1r 0005 mods1r 0006

02:40 execMODS mods.macsj0647_mos_1.obs mods1b 0003-0004 mods1r 0010-0011
  • the average seeing is 1.4 at the beginning of the integration
  • the average seeing for the second frame is 1.2

03:47 execMODS mods.macsj0647_mos_2.obs mods1b 0005-0006 mods1r 0012-0013
  • the average seeing is 1.0 at the beginning of the integration

04:54 execMODS mods.macsj0647_mos_3.obs mods1b 0007-0008 mods1r 0014-0015
  • the average seeing is 1.1 at the beginning of the integration
  • thin cirrus
  • 05:35 clouds coming in, b0008 and r0015 maybe not good

06:00 We slew to Gilli_4_Q1 J1048 Mask1
  • acqMODS 4Q1_recover.acq mods1r 0016-0018
  • sky almost clear now

06:18 execMODS 4Q1_J1048_M1_mos_6.obs mods1r 0019-0022
  • no clouds, seeing 1.3-1.5
  • in the last 15 minutes clouds

07:27 moving to Piconcelli_18_C3 RBS832
  • clouds, we are not guiding

07:39 slewing to Gilli_Q1, not guiding too much clouds

08:27 we try to acquire Piconcelli_18_C3 RBS832 mods1r 0025-0027, we do not see the target in the acquisition

08:52 we try to acquire Piconcelli_18_C3 RBS832 mods1r 0028-0029
  • new guiding star R=12.3
  • 08:59 modsAlign -l mod1r 0028 mods1r 0029
  • red readout counter stop updating, acquisition image 140204M8.00v.fits. After the last throught slit image, we commit a positive x-offset.

09:11 execMods C3_LS.obs mods1r 0032-0033
  • non detection due to thick clouds

09:35 the TO closes the dome: due to thick clouds he's not able to collimate

-- EleanoraSani - 05 Feb 2014
Topic revision: r7 - 14 Feb 2014, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback